OpenStudy (anonymous):

The difference between two numbers is 28. The larger number is 8 less than twice the smaller number. Find both numbers. Please help !

OpenStudy (anonymous):

x - y = 28 x = 2y - 8 now sub 2y - 8 in for x into the first equation 2y - 8 - y = 28 y = 28 + 8 y = 36 now sub 36 in for y into either of the equations to find x x - y = 28 x - 36 = 28 x = 28 + 36 x = 64 large number (x) = 64 smaller number (y) = 36

OpenStudy (anonymous):

Thank you !